### v1

# An Upper Bound on the Error Probability of RPA Decoding of Reed-Muller Codes Over the BSC

Abstract
In this paper, we revisit the Recursive Projection-Aggregation (RPA) decoder, of Ye and Abbe (2020), for Reed-Muller (RM) codes. Our main contribution is an explicit upper bound on the probability of incorrect decoding, using the RPA decoder, over a binary symmetric channel (BSC). Importantly, we focus on the events where a single iteration of the RPA decoder, in each recursive call, is sufficient for convergence. Key components of our analysis are explicit estimates of the probability of incorrect decoding of first-order RM codes using a maximum likelihood (ML) decoder, and estimates of the error probabilities during the aggregation phase of the RPA decoder. Our results allow us to show that for RM codes with blocklength $N=2^{m}$, the RPA decoder can achieve vanishing error probabilities, in the large blocklength limit, for RM orders that grow roughly logarithmically in $m$.
I Introduction
Reed-Muller (RM) codes are a family of binary linear codes that are obtained by the evaluations of Boolean polynomials on the points of the Boolean hypercube [1, 2]. They are of widespread interest in the information theory and error-control coding communities, owing to recent breakthrough theoretical progress has shown that RM codes by themselves are capacity-achieving for the binary erasure channel [3], and more generally, for BMS channels [4, 5]. Further, they share a close relationship [6, 7] with polar codes [8]—the first explicit family of (linear) codes that achieve capacity over binary-input memoryless symmetric (BMS) channels and also provably achieve high rates over many channels with memory [9, 10].
In the light of their importance, much work has been dedicated to finding good decoding algorithms for RM codes. The earliest such algorithm by Reed [1] is capable of correcting bit-flip errors up to half the minimum distance of the RM code. For the special case of first-order RM codes, a Fast Hadamard Transform-based (or FHT-based) decoder was designed in [11, 12], which is an efficient implementation of a maximum likelihood (ML) decoding procedure. For second-order RM codes, a decoding algorithm that works well for moderate blocklength $N$ ($N\leq 1024$) RM codes was discussed in [13, 14]. Further, recursive decoding algorithms for RM codes (with and without lists) with good performance at moderate blocklengths were described and analyzed in [15, 16, 17]. We also refer the reader to several recent data-driven and heuristic approaches for decoding RM codes [18, 19, 20, 21].
A much more recent decoding algorithm, variants of which were shown to achieve near-ML performance at moderate blocklengths over the binary symmetric channel (BSC), is the Recursive Projection-Aggregation (RPA) decoder of Ye and Abbe [22]. Via extensive simulation studies, the work [22] showed that the RPA algorithm outperforms the decoding algorithms in [13, 14] for second-order RM codes and the decoding algorithms in [15, 16, 17], for selected moderate blocklength RM codes. Later works [23, 24] presented procedures for reducing the complexity of the RPA algorithm. We mention however that recent work [25] suggests that running the RPA decoder by employing an aggregation step that uses only a subset of the total number of subspaces of a fixed dimension could lead to poor performance, due to the decoder’s potential failure in correcting so-called “coset error patterns".
While it is now well-accepted that the RPA algorithm (as described in [22]) demonstrates good performance in practice for low-rate (in particular, fixed-order) RM codes, a formal, theoretical analysis of the RPA decoder has been largely lacking in the literature. In this paper, we make progress towards this goal, by obtaining explicit upper bounds on the error probabilities of the RPA decoder, when the channel used is the BSC. Importantly, we prove that for sufficiently large blocklengths, with overwhelmingly high probability, the RPA decoder recovers the correct input codeword in RM codes of order that scales up to logarithmically in the parameter $m$, where $N=2^{m}$. This result can hence be viewed as supplementing the results in [26, 27], where it was shown that RM codes of order growing linearly in the parameter $m$, but with vanishing rate, have error probabilities that decay to zero in the large blocklength limit. Our result, although weaker in terms of the orders of the RM codes of interest, establishes that an explicit, low-complexity decoding algorithm (namely, the RPA algorithm) can in fact be used to achieve vanishing error probabilities.
Our analysis is based on explicit upper bounds on the probability of incorrect maximum likelihood (ML) decoding of first-order RM codes, which is used as a subroutine in the RPA decoder. With such upper bounds in place, we then characterize upper bounds on the probabilities of error in the “aggregation" steps carried out by the RPA decoder, conditioned on the event that all the “projections" are decoded correctly. Interestingly, our results show that even when one restricts attention to the event that each of the recursive calls to the RPA decoder converges in just one iteration, it is possible to obtain non-trivial performance estimates, which guarantee vanishing error probabilities for RM orders that increase with $m$.
An interesting line of future work will be to characterize the largest order of an RM code, as a function of the parameter $m$, which can be decoded via RPA decoding with vanishing error probability, in the limit as the blocklength increases to infinity. Such analysis will require a finer understanding of the correlations among the random variables that are recursed on over several iterations of the RPA decoder, when potentially several runs of each recursive call to the RPA decoder are run. Importantly, via such more sophisticated analysis, we shall be able to determine if RM codes can achieve positive rates over the BSC under RPA decoding.
II Notation and Preliminaries
II-A Notation
[…]
II-B Reed-Muller Codes
[…]
Definition II.1 (see Ch. 13 in [28], or [29]).
For $0\leq r\leq m$, the $r^{\text{th}}$-order binary Reed-Muller code RM$(m,r)$ is defined as
$\text{RM}(m,r):=\{\text{Eval}(f):f\in\mathbb{F}_{2}[x_{1},x_{2},\ldots,x_{m}],\ \text{deg}(f)\leq r\},$
where $\text{deg}(f)$ is the degree of the largest monomial in $f$, and the degree of a monomial $\prod_{j\in S:S\subseteq[m]}x_{j}$ is simply $|S|$.
It is known that the evaluation vectors of all the distinct monomials in the variables $x_{1},\ldots,x_{m}$ are linearly independent over $\mathbb{F}_{2}$. Hence, RM$(m,r)$ has dimension $\binom{m}{\leq r}:=\sum_{i=0}^{r}{m\choose i}$. We then have that RM$(m,r)$ is a $\left[2^{m},\binom{m}{\leq r}\right]$ linear code. It is also known that RM$(m,r)$ has minimum Hamming distance ${d}_{\text{min}}(\text{RM}(m,r))=2^{m-r}$.
II-C System Model and the RPA Decoder
In this subsection, we set up the communication system of interest and briefly recall the Recursive Projection-Aggregation (RPA) algorithm introduced in [22], for this setting.
II-C1 System Model
In this paper, we work with the setting of the transmission of a codeword of an RM code across a binary symmetric channel BSC$(p)$, where $p\in(0,\frac{1}{2})$ is the crossover probability of the channel. Let $\mathbf{c}$ be the codeword that is drawn from the code RM$(m,r)$, for $m\geq r\geq 1$ (the case $r=0$ corresponds to a simple repetition code), for transmission. We write $N:=2^{m}$ as the blocklength of the code. The codeword $\mathbf{c}$ is transmitted across a BSC$(p)$, resulting in the (random) received vector $\mathbf{Y}$, where
$\mathbf{Y}=\mathbf{c}\oplus\boldsymbol{\nu},$
where $\boldsymbol{\nu}\sim\text{Ber}^{\otimes N}(p)$. We index the coordinates of $\mathbf{Y}$ using vectors $\mathbf{z}\in\{0,1\}^{m}$ in lexicographic order. The RPA decoder will use this received vector $\mathbf{Y}$ to obtain an estimate of the transmitted codeword $\mathbf{c}$.
II-C2 The RPA Decoder
For any one-dimensional subspace $\mathbb{B}_{i}$ of $\{0,1\}^{m}$, $i\in[N-1]$, that contains the non-zero vector $\mathbf{b}_{i}$, we let $\{0,1\}^{m}/\mathbb{B}_{i}$ denote the collection of cosets of $\mathbb{B}_{i}$ in $\{0,1\}^{m}$. For any coset $T\in\{0,1\}^{m}/\mathbb{B}_{i}$ that contains the vector $\mathbf{z}$ (we sometimes write such a coset as $[\mathbf{z}+\mathbb{B}_{i}]$) and for the given received vector $\mathbf{Y}$, we define $\mathbf{Y}_{/\mathbb{B}_{i}}(T):=\mathbf{Y}_{\mathbf{z}}\oplus\mathbf{Y}_{\mathbf{z}\oplus\mathbf{b}_{i}}$ as the “projection" of the vector $\mathbf{Y}$ onto the coset $T$. Finally, we define the projection of $\mathbf{Y}$ onto the cosets of $\mathbb{B}_{i}$ as
$\mathbf{Y}_{/\mathbb{B}_{i}}:=\left(\mathbf{Y}_{/\mathbb{B}_{i}}(T):\ T\in\{0,1\}^{m}/\mathbb{B}_{i}\right),$ (1)
for some fixed ordering among cosets $T$. One can define a projection $\mathbf{c}_{/\mathbb{B}_{i}}$ analogously for the input codeword $\mathbf{c}$ too; via standard properties of Reed-Muller codes, it can be argued that $\mathbf{c}_{/\mathbb{B}_{i}}\in\text{RM}(m-1,r-1)$, if $\mathbf{c}\in\text{RM}(m,r)$. We now briefly recall the RPA decoder for second-order RM codes. The main decoding algorithm is shown as Algorithm 1, which in turn uses the “aggregation" step in Algorithm 2 as a subroutine. We observe that the RPA decoding procedure consists of two key steps: (i) decoding of first-order ($r=1$) RM codes via the Fast Hadamard Transform (FHT) decoder (see [12] and [29, Sec. V.B.1]), which is simply an efficient implementation of a maximum likelihood (ML) decoder for first-order RM codes, and (ii) an aggregation step that aggregates multiple noisy estimates of a single codeword symbol.
We mention that our presentation of the RPA decoding algorithm and the aggregation subroutine are specialized to the setting where one projects of the received vector $\mathbf{Y}$ using one-dimensional subspaces, as against the more general formulation in [22], which allows for the use of subspaces of higher dimension. As we shall see, projections via one-dimensional subspaces by themselves give rise to correct decoding, with very high probability, for sufficiently large blocklengths $N$.
Algorithm 1 The RPA decoding algorithm for RM$(m,r)$
1:procedure RPA($\mathbf{Y}$, $m$, $r$, $M_{\text{iter}}$)
2:     Set $N\leftarrow 2^{m}$.
3:     for $j=1:M_{\text{iter}}$ do
4:         Compute the projections $\mathbf{Y}_{/\mathbb{B}_{i}}$, $i\in[N-1]$ as in (1).
5:         if $r\neq 2$ then
6:              For each $i\in[N-1]$, set $\widehat{\mathbf{Y}}_{/\mathbb{B}_{i}}\leftarrow$ RPA($\mathbf{Y}_{/\mathbb{B}_{i}},m-1,r-1,M_{\text{iter}}$).
7:         else
8:              For each $i\in[N-1]$, decode $\mathbf{Y}_{/\mathbb{B}_{i}}$ to $\widehat{\mathbf{Y}}_{/\mathbb{B}_{i}}$ using the FHT-based decoder.
9:         Compute $\overline{\mathbf{Y}}\leftarrow$ Aggregate$(\mathbf{Y},\widehat{\mathbf{Y}}_{/\mathbb{B}_{1}},\ldots,\widehat{\mathbf{Y}}_{/\mathbb{B}_{N-1}})$.
10:         if $\mathbf{Y}=\overline{\mathbf{Y}}$ then break
11:         Update $\mathbf{Y}\leftarrow\overline{\mathbf{Y}}$.
12:     Return $\mathbf{Y}$.
Algorithm 2 The aggregation subroutine
1:procedure Aggregate($\mathbf{Y},\widehat{\mathbf{Y}}_{/\mathbb{B}_{1}},\ldots,\widehat{\mathbf{Y}}_{/\mathbb{B}_{N-1}}$)
2:     Initialize the length-$N$ vector $\textsc{Flip}\leftarrow\mathbf{0}$.
3:     for $\mathbf{x}\in\{0,1\}^{m}$ do
4:         Compute $\phi(\mathbf{x})=\sum_{i=1}^{N-1}\mathds{1}\{Y_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\neq\widehat{Y}_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\}$.
5:         Set Flip$(\mathbf{x})=1$, if $\phi(\mathbf{x})>\frac{N-1}{2}$.
6:     Set $\overline{Y}\leftarrow\mathbf{Y}\oplus\textsc{Flip}$.
7:     Return $\overline{Y}$.
III Main Results
Given a fixed Reed-Muller code RM$(m,r)$, suppose that the (fixed) codeword $\mathbf{c}\in\text{RM}(m,r)$ is transmitted across the BSC$(p)$, and let $\widehat{\textbf{C}}$ denote the estimate produced by the RPA decoder. We then define the probability of error on transmission of $\mathbf{c}$ as $P_{\text{err}}(\mathbf{c}):=\Pr\left[\widehat{\mathbf{C}}\neq\mathbf{c}\mid\mathbf{c}\ \text{transmitted}\right]$ and the probability of error of the code RM$(m,r)$, under RPA decoding, when a codeword $\mathbf{C}$ is chosen uniformly at random from the code as $P_{\text{err}}(\text{RM}(m,r)):=\Pr\left[\widehat{\mathbf{C}}\neq\mathbf{C}\right]$.
An important fact about the error probabilities above is [22, Prop. 2], which states that for all $m\geq 1,1\leq r\leq m$, we have that the probability of error over the BSC$(p)$, upon transmission of a codeword $\mathbf{c}\in\text{RM}(m,r)$ is independent of $\mathbf{c}$. This, in turn, implies that we can focus our analysis on the case when $\mathbf{c}=\mathbf{0}\in\{0,1\}^{N}$, with $P_{\text{err}}(\text{RM}(m,r))=P_{\text{err}}(\mathbf{0})$. In all the analysis that follows, we shall make the simplifying assumption that the input codeword is fixed to be the all-zeros codeword $\mathbf{0}$.
Before we state our main result, we will require some more notation. Let $\overline{p}:=\frac{1}{2}\cdot(1-(1-2p)^{2^{r-2}})$ and for any $\alpha\in(0,\frac{1}{2})$, let $\eta(\alpha):=\frac{1}{2}\cdot(1-4\alpha(1-\alpha))$. Note that $\eta(\alpha)<\frac{1}{2}$, for all $\alpha\in(0,\frac{1}{2})$. Our main result is encapsulated in the following theorem:
Theorem III.1.
For any $0<\epsilon<\eta(\overline{p})$, we have that for $r\geq 2$,
$P_{\text{err}}(\text{RM}(m,r))\leq 32N^{r+1}\cdot e^{-2^{-(r+1)}\cdot{N\epsilon^{2}}}.$
It can easily be checked that for any fixed (constant) $r\geq 2$, we have that
$\lim_{m\to\infty}P_{\text{err}}(\text{RM}(m,r))=\lim_{N\to\infty}P_{\text{err}}(\text{RM}(m,r))=0.$
In fact, the following stronger corollary holds, by simply evaluating the limit of the right-hand side in Theorem III.1. Let $c=c(p):=\frac{\log 2}{\log\left(\frac{1}{1-2p}\right)}$.
Corollary III.1.
For any $0<\overline{c}<c$, we have that for all $r\leq\log(\overline{c}m)$,
$\lim_{m\to\infty}P_{\text{err}}(\text{RM}(m,r))=0.$
Proof.
First, observe that
$\eta(\overline{p})$ $=\frac{1}{2}\cdot\left(1-4\overline{p}(1-\overline{p})\right)$
$=\frac{1}{2}\cdot\left(1-\left(1-(1-2p)^{2^{r-2}}\right)\cdot\left(1+(1-2p)^{2^{r-2}}\right)\right)$
$=\frac{1}{2}\cdot\left(1-2p\right)^{2^{r-1}}.$
Now, from the upper bound in Theorem III.1, for any fixed $\delta\in(0,1)$, writing $\epsilon^{2}=\delta\cdot\eta(\overline{p})^{2}$, we have $P_{\text{err}}(\text{RM}(m,r))\leq 2^{\rho_{m}(r,\delta)}$, where
$\rho_{m}(r,\epsilon):=m(r+1)-\frac{\delta\log_{2}(e)}{8}\cdot 2^{m-r}\cdot(1-2p)^{2^{r}}+5.$
By picking $r\leq\log(\overline{c}m)$, for any $\overline{c}<c$, it can be checked that the corollary indeed holds. ∎
Note, however, that RM codes with $r\leq\log(cm)$ as above have rate $R_{m}:=\frac{\log|\text{RM}(m,r)|}{N}=\frac{{m\choose\leq r}}{N}$ such that $R_{m}\xrightarrow{m\to\infty}0$. In other words, via the results in this paper, we only obtain that RPA decoding achieves vanishing error probabilities, in the limit as the blocklength goes to infinity, for selected RM codes that have vanishing rate. An important question hence, as mentioned earlier, is whether RM codes can achieve positive rate (or even the capacity of the BSC$(p)$) under RPA decoding—a fertile topic for future research.
Furthermore, for fixed (constant) orders $r$, Theorem III.1 also allows us to understand how quickly, as a function of $m$, the crossover probability $p$ (also called the “decoding radius") can converge to its maximum value that is $1/2$, while also ensuring that $P_{\text{err}}(\text{RM}(m,r))\xrightarrow{m\to\infty}0$. Observe that the statement that a decoding algorithm can achieve vanishing error probabilities, in the limit as the blocklength $N$ tends to infinity, is equivalent to the statement that the algorithm can correct $N(p-\delta)$ errors with high probability, for any $\delta\in(0,1)$. The following result hence establishes a lower bound on the number of such errors correctable with high probability, as a function of $m$, for fixed orders $r$, in a vein similar to the results in [15].
Corollary III.2.
For fixed $r\geq 2$ and any $\delta,\beta\in(0,1)$, for sufficiently large $m$, the RPA decoder can correct $\frac{N}{2}\cdot\left(1-\gamma_{m}-\delta\right)$ errors with high probability, where
$\gamma_{m}=\left(\frac{8(5+m(r+1+\beta))}{d\cdot\log_{2}e}\right)^{\frac{1}{2^{r}}}.$
Proof.
Fix $r\geq 2$. In order to ensure that $P_{\text{err}}(\text{RM}(m,r))\xrightarrow{m\to\infty}0$, from Theorem III.1, we would like $p$ to be such that for some fixed $\beta\in(0,1)$,
$e^{2^{-(r+1)}\cdot N\eta(\overline{p})^{2}}\geq 32N^{r+1+\beta}.$ (2)
By using the fact that $\eta(\overline{p})=\frac{1}{2}\cdot(1-2p)^{2^{r-1}}$ (see the proof of Corollary III.1), we see via straightforward algebraic manipulations that (2) is equivalent to requiring that $p\leq\frac{1}{2}\cdot\left(1-\gamma_{m}\right),$ where $\gamma_{m}$ is as in the statement of the corollary. We hence obtain that for sufficiently large $m$, the RPA decoder can correct $N(p-\delta)=\frac{N}{2}\cdot\left(1-\gamma_{m}-\delta\right)$ errors with high probability, for any fixed $\delta\in(0,1)$. ∎
The result in Corollary III.2 allows for a direct comparison with the main results in [15]. In particular, [15, Thm. 2] shows that there exists a decoder for RM$(m,r)$, for fixed orders $r$, which can correct $\frac{N}{2}\cdot\left(1-\beta_{m}\right)$ errors with high probability, for sufficiently large $m$, where $\beta_{m}=\left(\frac{qm}{d}\right)^{\frac{1}{2^{r}}}$, for $q>\ln 4$. Note that for $r\geq 2$, we have
$\gamma_{m}\geq\left(\frac{24m}{d\cdot\log_{2}e}\right)^{\frac{1}{2^{r}}}>\left(\frac{2m\cdot\ln 4}{d}\right)^{\frac{1}{2^{r}}}.$
This implies that the convergence of the decoding radius to $1/2$ (captured by $\gamma_{m}$), for fixed orders $r\geq 2$, under our analysis of RPA decoding, is slower than that obtainable using the decoding algorithm $\Phi_{r}^{m}$ in [15] (captured by $\beta_{m}$). However, for sufficiently large $m$, it can be verified similarly that the convergence of the decoding radius to $1/2$ under RPA decoding is faster than the decoding algorithm $\Psi_{r}^{m}$ in [15]. It is an interesting question to check whether the RPA decoder can outperform $\Phi_{r}^{m}$ in [15] in terms of the speed of convergence of the decoding radius, using more sophisticated analysis of the RPA error probabilities, possibly by considering cases where each RPA run converges after larger than one iteration, or by using higher dimensional subspaces for projection.
In Sections IV and V, we make progress towards proving III.1, formally. In the next section, we shall analyze the behaviour of the FHT decoder in Step 8 of Algorithm 1 above (which acts on the first-order RM codes obtained via the projection step) and the aggregation step in Step 9 of Algorithm 1, separately. Via such analysis, we shall obtain an understanding of the performance of the RPA decoder on second-order ($r=2$) RM codes. In Section V, we shall use the results obtained in the next section, recursively, to get estimates of the error probabilities of the RPA decoder for general orders $r>2$.
IV Analysis on Second-Order RM Codes
In this section, we investigate the performance of the RPA decoder on second-order RM codes, i.e., codes of the form RM$(m,2)$, $m\geq 2$, which are used for transmission over a BSC$(p)$. In particular, our interest is in establishing bounds on the probability that the (iterative) RPA decoder converges to the true transmitted codeword, which is assumed to be drawn uniformly from a second-order RM code.
Let $\mathbf{Y}\in\{0,1\}^{N}$ denote the (random) vector received by passing the input codeword $\mathbf{0}\in\text{RM}(m,2)$ through the BSC$(p)$. We hence have that $\mathbf{Y}\sim\text{Ber}^{\otimes{N}}(p)$. Further, by the definition of the (random) vectors $\mathbf{Y}_{/\mathbb{B}_{i}}$, for any $i\in[N-1]$, it follows that $\mathbf{Y}_{/\mathbb{B}_{i}}\sim\text{Ber}^{\otimes(N/2)}(2p(1-p))$.
IV-A The Performance of the FHT Decoder
Recall from Section II-C that the FHT-based decoder for first-order RM codes is simply an efficient implementation of ML decoding. In this section, we hence directly analyze the behaviour of the ML decoder on first-order RM codes, via a slightly different perspective that is closely related to, but different in form from, the FHT-based decoder.
Now, fix a one-dimensional subspace $\mathbb{B}_{i}$, for some $i\in[N-1]$, as in Algorithm 1; let $\mathbf{b}_{i}\in\{0,1\}^{N}$ denote the non-zero vector in $\mathbb{B}_{i}$. Given the received vector $\mathbf{Y}$, we have that $\mathbf{Y}_{/\mathbb{B}_{i}}\in\text{RM}(m-1,1)$. Consider the ML (or FHT-based) decoder in Step 8 of Algorithm 1 that acts on $\mathbf{Y}_{/\mathbb{B}_{i}}$. First, we transform $\mathbf{Y}_{/\mathbb{B}_{i}}\in\{0,1\}^{N/2}$ to the vector $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}\in\{-1,1\}^{N/2}$, via the standard mapping in Section II-A.
Further, consider the family of functions $\left(\chi_{\mathbf{s}}:\mathbf{s}\in\{0,1\}^{m-1}\right)$, where $\chi_{\mathbf{s}}(\mathbf{x}):=(-1)^{\mathbf{x}\cdot\mathbf{s}}$, $\mathbf{x}\in\{0,1\}^{m-1}$. We interchangably treat each $\chi_{\mathbf{s}}$ (and likewise, $\mathbf{Y}_{/\mathbb{B}_{i}}$) as a function from $\{0,1\}^{m-1}$ to $\{-1,1\}$, and as a length-$N/2$ vector over $\{-1,1\}$. It can easily be checked that there exists a one-one correspondence between codewords of RM$(m-1,1)$ and the collection of vectors $\chi:=\left(\chi_{\mathbf{s}}:\mathbf{s}\in\{0,1\}^{m-1}\right)\cup\left(-\chi_{\mathbf{s}}:\mathbf{s}\in\{0,1\}^{m-1}\right)$. Note that the all-zeros codeword $\mathbf{0}\in\text{RM}(m-1,1)$ corresponds to the vector $+\chi_{\mathbf{0}}$. Thus, the ML decoder ML for the first-order RM$(m-1,1)$ code, given $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}$, essentially performs:
$\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})=\operatorname*{arg\,min}\limits_{f\in\chi}\ \text{dist}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},f),$ (3)
where for functions $f,g:\{0,1\}^{n}\to\{-1,1\}$, we define
$\text{dist}(f,g):=\frac{1}{2^{n}}\cdot\sum_{\mathbf{x}\in\{0,1\}^{n}}\mathds{1}\{f(\mathbf{x})\neq g(\mathbf{x})\}.$
Further, given $f,g:\{0,1\}^{n}\to\{-1,1\}$, we define their inner product
$\langle f,g\rangle:=\frac{1}{2^{n}}\cdot\sum_{\mathbf{x}\in\{0,1\}^{n}}f(\mathbf{x})g(\mathbf{x}).$
Via simple manipulations (see, e.g., [30, Prop. 1.9]), we have the following lemma:
Lemma IV.1.
For functions $f,g:\{0,1\}^{n}\to\{-1,1\}$, we have
$\langle f,g\rangle=1-2\text{dist}(f,g).$
Thus, the task in (3) reduces to:
$\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})$ $=\operatorname*{arg\,max}\limits_{f\in\chi}\ \langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},f\rangle,$
which, in turn, identifies $\mathbf{s}\in\{0,1\}^{m-1}$ and $\sigma\in\{-1,1\}$ that maximize $\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\sigma\cdot\chi_{\mathbf{s}}\rangle.$ We hence write
$\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})=\operatorname*{arg\,max}\limits_{\mathbf{s}\in\{0,1\}^{m-1},\ \sigma\in\{-1,1\}}\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\sigma\cdot\chi_{\mathbf{s}}\rangle.$ (4)
As mentioned earlier, the all-zeros codeword corresponds to $\mathbf{s}=\mathbf{0}$ and $\sigma=1$. We use the expression (4) above as the structure of our ML decoder, in what follows. We mention here that for a function $f:\{0,1\}^{n}\to\{-1,1\}$, the term $\langle f,\chi_{\mathbf{s}}\rangle$ is precisely the Fourier coefficient $\widehat{f}(\mathbf{s})$ (see [30] for more details on the Fourier analysis of Boolean functions) of $f$ at the vector $\mathbf{s}\in\{0,1\}^{m-1}$.
The following key theorem shows that with overwhelming probability, for sufficiently large $N$, the ML decoder, when acting on $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}$, produces the correct input codeword $\mathbf{c}\in\text{RM}(m-1,1)$, which is the all-zeros codeword. Let $\eta=\eta(p):=\frac{1}{2}\cdot(1-4p(1-p))$.
Theorem IV.1.
We have that for any $\epsilon<\eta(p)$,
$\Pr[\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})=(\mathbf{0},1)]\geq 1-8N\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
The proof of Theorem IV.1 makes use of the following simple helper lemmas.
Lemma IV.2.
For all $\epsilon>0$ and any $i\in[N-1]$, we have that
$\Pr\left[\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{0}}\rangle-(1-4p(1-p))\rvert\leq\epsilon\right]\geq 1-2e^{-\frac{N\epsilon^{2}}{4}}.$
Proof.
Note that
$\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{0}}\rangle$ $=\frac{1}{{N/2}}\cdot\sum_{\mathbf{x}\in\{0,1\}^{m-1}}\left({Y}^{\pm}_{/\mathbb{B}_{i}}\right)_{\mathbf{x}}$
$=\frac{2}{N}\cdot\sum_{\mathbf{x}\in\{0,1\}^{m-1}}Z_{\mathbf{x}},$
where $Z_{\mathbf{x}}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{Z}$, where $P_{Z}$ is a Rademacher distribution with $P_{Z}(-1)=1-P_{Z}(1)=2p(1-p)$. Clearly, we have that $\mathbb{E}\left[\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{0}}\rangle\right]=1-4p(1-p)$. The claim then follows as a direct consequence of Hoeffding’s inequality [31] (see also [32, Thm. 2.2.6]). ∎
Lemma IV.3.
For all $\epsilon>0$ and any $i\in[N-1]$, we have that for any $\mathbf{s}\neq\mathbf{0}$,
$\Pr\left[\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle\rvert\leq\epsilon\right]\geq 1-4e^{-\frac{N\epsilon^{2}}{8}}.$
Proof.
Using the fact that $\mathbf{Y}_{/\mathbb{B}_{i}}\sim\text{Ber}^{\otimes(N/2)}(2p(1-p))$, we see that
$\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle$ $=\frac{2}{N}\cdot\sum_{\mathbf{x}\in\{0,1\}^{m-1}}\left({Y}^{\pm}_{/\mathbb{B}_{i}}\right)_{\mathbf{x}}$
$=\frac{2}{N}\cdot\left[\sum_{j=1}^{N/4}Z^{\prime}_{j}-\sum_{k=1}^{N/4}Z^{\prime\prime}_{k}\right],$
where $Z^{\prime}_{j},Z^{\prime\prime}_{k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{Z}$, where $P_{Z}$ is a Rademacher distribution with $P_{Z}(-1)=1-P_{Z}(1)=2p(1-p)$. The second equality above holds since the vector $\chi_{\mathbf{s}}$, for $\mathbf{s}\neq\mathbf{0}$ has exactly $N/4$ coordinates that are $-1$ and $N/4$ coordinates that are $+1$. Let $\alpha_{1}:=\frac{2}{N}\cdot\sum_{j=1}^{N/4}Z^{\prime}_{j}$ and $\alpha_{2}:=\frac{2}{N}\cdot\sum_{j=1}^{N/4}Z^{\prime\prime}_{j}$. Note that $\mathbb{E}[\alpha_{1}]=\mathbb{E}[\alpha_{2}]=\frac{1}{2}\cdot(1-4p(1-p)):=\beta$. Now, we have that
$\Pr\left[\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle\rvert\leq\epsilon\right]$
$=\Pr[|\alpha_{1}-\alpha_{2}|\leq\epsilon]$
$\geq\Pr\left[\alpha_{1}\in[\beta-\epsilon/2,\beta+\epsilon/2]\text{ and }\alpha_{2}\in[\beta-\epsilon/2,\beta+\epsilon/2]\right]$
$\geq\left(1-2e^{-\frac{N\epsilon^{2}}{8}}\right)^{2}\geq 1-4e^{-\frac{N\epsilon^{2}}{8}},$
where the second inequality holds via an application of Hoeffding’s inequality [31] (see also [32, Thm. 2.2.6]). ∎
The proof of Theorem IV.1 then follows easily by combining Lemmas IV.2 and IV.3 above.
Proof of Theorem IV.1.
Fix any $0<\epsilon<\eta(p)$. First, observe that by a union bound argument, the intersection of the events
$\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{0}}\rangle-(1-4p(1-p))\rvert\leq\epsilon$
and $\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle\rvert\leq\epsilon$, for all $\mathbf{s}\neq\mathbf{0}$, occurs with probability at least $1-8N\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ In other words, we have that with probability at least $1-8N\cdot e^{-\frac{N\epsilon^{2}}{8}}$, the following events occur: $\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{0}}\rangle\in[(1-4p(1-p))-\epsilon,(1-4p(1-p))+\epsilon]$ and $\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},-\chi_{\mathbf{0}}\rangle\in[-(1-4p(1-p))-\epsilon,-(1-4p(1-p))+\epsilon]$; furthermore, for all $\mathbf{s}\neq\mathbf{0}$, we have $\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle\in[-\epsilon,\epsilon]$ and $\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},-\chi_{\mathbf{s}}\rangle\in[-\epsilon,\epsilon]$. From our choice of $\epsilon<\eta(p)$, it can thus be verified that with probability at least $1-8N\cdot e^{-\frac{N\epsilon^{2}}{8}}$, we have
$\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{0}}\rangle=\max_{f\in\chi}\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},f\rangle.$
thereby concluding the proof of the theorem. ∎
With Theorem IV.1 in place, a simple application of the union bound then results in the following result:
Proposition IV.1.
We have that for any $\epsilon<\eta(p)$,
$\Pr\left[\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})=(\mathbf{0},1),\text{for all }i\in[N-1]\right]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
In other words, we have that for any $\epsilon<\eta(p)$,
$\Pr\left[\widehat{Y}_{/\mathbb{B}_{i}}=\mathbf{0},\text{for all }i\in[N-1]\right]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
We shall use this fact in the next subsection to argue that with very high probability, for sufficiently large $N$, the aggregation step produces the correct, all-zeros input codeword in just a single iteration of the RPA decoder.
IV-B The Aggregation Step
Let $\mathcal{G}$ denote the event $\left\{\widehat{Y}_{/\mathbb{B}_{i}}=\mathbf{0},\text{for all }i\in[N-1]\right\}$. Proposition IV.1 shows that $\Pr[\mathcal{G}]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ In this subsection, we analyze the performance of the aggregation step, conditioned on the event $\mathcal{G}$. Observe that, conditioned on $\mathcal{G}$, the aggregation step computes, for each $\mathbf{x}\in\{0,1\}^{m}$, the function $\phi(\mathbf{x})=:\phi^{(N)}(\mathbf{x})=\sum_{i=1}^{N-1}\mathds{1}\{Y_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\neq\widehat{Y}_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\}$. Upon conditioning on $\mathcal{G}$, it can be checked that
$\phi(\mathbf{x})=\sum_{i=1}^{N-1}\left(Y_{\mathbf{x}}\oplus Y_{\mathbf{x}+\mathbf{b}_{i}}\right),$
where the summation over $i\in[N-1]$ above is over the reals. The aggregation step then sets Flip$(\mathbf{x})=\textsf{Flip}^{(N)}(\mathbf{x})=1$, if $\phi(\mathbf{x})>\frac{N-1}{2}$, and Flip$(\mathbf{x})=0$, otherwise. Finally, the vector $\overline{\mathbf{Y}}=\mathbf{Y}\oplus\textsf{Flip}$ is computed, which then serves as the new “received vector" in the next iteration.
In what follows, we argue that with high probability, for large enough $N$, the vector Flip equals the vector $\mathbf{Y}$, thereby leading to $\overline{\mathbf{Y}}=\mathbf{0}$. It can then be checked that once $\overline{\textbf{Y}}$ equals the all-zeros vector at the end of the first iteration of the RPA decoder, further iterations only produce the all-zeros vector, implying that the RPA decoder has converged to the correct input codeword after just one iteration.
Now, observe that
$\phi(\mathbf{x})=\begin{cases}\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}},\ \text{if $Y_{\mathbf{x}}=0$},\\ N-1-\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}},\ \text{if $Y_{\mathbf{x}}=1$}.\end{cases}$
where each $Y_{\mathbf{z}}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Ber}(p)$. We define $\overline{\phi}(\mathbf{x})=\overline{\phi}^{(N)}(\mathbf{x})=\frac{\phi(\mathbf{x})}{N-1}$ and $\overline{\phi}_{\infty}(\mathbf{x}):=p(1-Y_{\mathbf{x}})+(1-p)Y_{\mathbf{x}}$. It is not too hard to believe that for large enough $N$, conditioned on the event $\mathcal{G}$, the random variable $\textsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\}$ is close to $\mathds{1}\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\}$. The following proposition thus holds:
Proposition IV.2.
For all $\epsilon<\eta(p)$, we have that for any $\mathbf{x}\in\{0,1\}^{m}$,
$\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\bigg{|}\mathcal{G}\right]\geq$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 1-16N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
Before we prove the above proposition, we state and prove a useful lemma:
Lemma IV.4.
For any $\epsilon>0$, we have that for any $\mathbf{x}\in\{0,1\}^{m}$,
$\Pr\left[\left\lvert\overline{\phi}^{(N)}(\mathbf{x})-\overline{\phi}_{\infty}(\mathbf{x})\right\rvert\leq\epsilon\right]\geq 1-2\cdot e^{-2(N-1)\epsilon^{2}}.$
Proof.
First, observe that $\mathbb{E}\left[Y_{\mathbf{z}}\right]=p$, for any $\mathbf{z}\in\{0,1\}^{m}$. Further, we can write, for any $\mathbf{x}\in\{0,1\}^{m}$,
$\overline{\phi}^{(N)}(\mathbf{x})$ $=\left(\frac{1}{N-1}\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}}\right)\cdot(1-Y_{\mathbf{x}})+Y_{\mathbf{x}}\cdot\left(1-\frac{1}{N-1}\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}}\right).$ (5)
Now, via Hoeffding’s inequality [31] (see also [32, Thm. 2.2.6]), with probability at least $1-2\cdot e^{-2(N-1)\epsilon^{2}}$, we have
$\left\lvert\frac{1}{N-1}\cdot\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}}-p\right\rvert\leq\epsilon.$
Substituting this back in (5), we obtain that with probability at least $1-2\cdot e^{-2(N-1)\epsilon^{2}}$, we must have $\left\lvert\overline{\phi}^{(N)}(\mathbf{x})-\overline{\phi}_{\infty}(\mathbf{x})\right\rvert\leq\epsilon,$ thereby proving the lemma. ∎
While the lemma above shows that with high probability, the random variables inside the indicator functions in $\mathsf{Flip}^{(N)}(\mathbf{x})$ and $\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}$ are close, for large enough $N$, Proposition IV.2 claims that these indicator functions are themselves close, with high probability, when conditioned on the event $\mathcal{G}$. We now prove Proposition IV.2.
Before we proceed, we introduce one additional piece of notation: let $\overline{\mathcal{G}}$ denote the event $\left\{\left\lvert\overline{\phi}^{(N)}(\mathbf{x})-\overline{\phi}_{\infty}(\mathbf{x})\right\rvert\leq\epsilon\right\}$.
Proof of Proposition IV.2.
We shall first prove that unconditionally, for any $\mathbf{x}\in\{0,1\}^{m}$, the random variable $\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}$ is close to the random variable $\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}$, with high probability, if $N$ is sufficiently large. This will then allow us to deduce the result of the proposition, by noting that conditioned on the event $\mathcal{G}$, we have that $\textsc{Flip}^{(N)}(\mathbf{x})$ equals $\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}$.
To this end, observe that the following sequence of inequalities holds:
$\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right]$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\ \bigg{|}\ \overline{\mathcal{G}}\right]\cdot\Pr[\overline{\mathcal{G}}]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\ \bigg{|}\ \overline{\mathcal{G}}\right]\cdot(1-\delta),$
where $\delta:=2e^{-2(N-1)\epsilon^{2}}.$ Here, inequality (a) holds via the law of total probability, and (b) holds via Lemma IV.4. We now claim that conditioned on $\overline{\mathcal{G}}$, the event
$\mathcal{A}^{\prime}:=\left\{\left\lvert\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})-\epsilon>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right\}$
implies the event
$\mathcal{A}:=\left\{\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right\}.$
To this end, it suffices to show that the expression $g_{1}(\mathbf{x}):=\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert$ is at most the expression $g_{2}(\mathbf{x}):=\left\lvert\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})-\epsilon>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert$, if $\overline{\mathcal{G}}$ holds. Indeed, observe that since $\epsilon<1/2$, we have from the event $\overline{\mathcal{G}}$ that
$\overline{\phi}_{\infty}(\mathbf{x})-\frac{1}{2}\geq\overline{\phi}(\mathbf{x})-\frac{1}{2}\geq\overline{\phi}_{\infty}(\mathbf{x})-\epsilon-\frac{1}{2},$
thereby leading to
$\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\leq\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}\leq\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})-\epsilon>\frac{1}{2}\right\}.$
Hence, we have that
$g_{1}(\mathbf{x})$ $=\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}$
$\leq\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})-\epsilon>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}=g_{2}(\mathbf{x}),$
thereby proving the claim.
Hence, carrying on from (b) above, we get that since the event $\mathcal{A}^{\prime}$ implies the event $\mathcal{A}$, conditioned on $\overline{\mathcal{G}}$,
$\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right]$
$\geq\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})-\epsilon>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right]\cdot(1-\delta)$
$=\Pr\left[\left\lvert\mathds{1}\left\{Y_{\mathbf{x}}>\frac{\frac{1}{2}+\epsilon-p}{1-2p}\right\}-\mathds{1}\left\{Y_{\mathbf{x}}>\frac{\frac{1}{2}-p}{1-2p}\right\}\right\rvert\leq\epsilon\right]\cdot(1-\delta)$
$=\Pr\left[\mathds{1}\left\{Y_{\mathbf{x}}>\frac{\frac{1}{2}+\epsilon-p}{1-2p}\right\}=\mathds{1}\left\{Y_{\mathbf{x}}>\frac{1}{2}\right\}\right]\cdot(1-\delta)$
$=\left(\Pr\left[Y_{\mathbf{x}}<\frac{1}{2}\right]+\Pr\left[Y_{\mathbf{x}}>\frac{\frac{1}{2}+\epsilon-p}{1-2p}\right]\right)\cdot(1-\delta)=1-\delta,$
since $0<\epsilon<\eta(p)\leq\frac{1}{2}-p$, for all $p<1/2$.
Now, let $\delta^{\prime}:=8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}$. The following inequalities then hold:
$\Pr\left[\left\lvert\mathsf{Flip}^{(N)}(\mathbf{x})-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\ \bigg{|}\ \mathcal{G}\right]$
$=\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\ \bigg{|}\ \mathcal{G}\right]$
$\geq{\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right]-\Pr[\mathcal{G}^{c}]}$
$\geq 1-\delta-\delta^{\prime}\geq 1-2\delta^{\prime}.$
Here, we define $\mathcal{G}^{c}:=\left\{\widehat{Y}_{/\mathbb{B}_{i}}\neq\mathbf{0},\text{for some }i\in[N-1]\right\}$, which leads to the first inequality above by simple manipulations of the law of total probability. The second inequality above then makes use of Proposition IV.1.
The proof of the proposition then follows by observing that $\mathsf{Flip}^{(N)}(\mathbf{x})$ and $\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}$ are both Bernoulli random variables, and hence their closeness must imply their equality, since $\epsilon<\eta(p)<1$. ∎
Thus, after some manipulation, we have obtained, for any $\mathbf{x}\in\{0,1\}^{m}$, conditioned on the event $\mathcal{G}$, a simple expression for the structure of the random variable $\mathsf{Flip}^{(N)}(\mathbf{x})$ that holds for sufficiently large $N$, with high probability. All that remains to be shown is that the random vector $\mathsf{Flip}^{(N)}$, with high probability, equals the received vector $\mathbf{Y}$. The following theorem shows that this fact is indeed true:
Theorem IV.2.
We have that for all $\epsilon<\eta(p)$,
$\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\right]\geq 1-32N^{3}\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
Proof.
First, observe that by a union bound argument, we have from Proposition IV.2 that
$\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\},\text{ for all $\mathbf{x}$}\ \bigg{|}\ \mathcal{G}\right]$
$\geq 1-16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ (6)
Next, we note that the random variable $\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}$ equals $\mathds{1}\left\{Y_{\mathbf{x}}>\frac{1}{2}\right\}$, which, in turn, equals $Y_{\mathbf{x}}$ itself. Hence, (6) above simply says that
$\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\ \bigg{|}\ \mathcal{G}\right]\geq 1-16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
Finally, note that the unconditional probability
$\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\right]$ (7)
$\geq\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\ \bigg{|}\ \mathcal{G}\right]\cdot\Pr[\mathcal{G}]$ (8)
$\geq\left(1-16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}\right)\left(1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}\right)$
$\geq 1-16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}$
$\geq 1-32N^{3}\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ (9)
Here, the second inequality above holds via Proposition IV.1. ∎
Thus, we finally obtain the following corollary, which was the statement we were aiming to prove:
Figure 1: Figure representing a projection-aggregation tree
Corollary IV.1.
After a single iteration of the RPA decoder for second-order RM codes, we have for all $\epsilon<\eta(p)$ that
$\Pr[\overline{\mathbf{Y}}=\mathbf{0}]\geq 1-32N^{3}\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
Thus, we see that with overwhelming probability, given any codeword from a second-order RM code that is transmitted across a BSC$(p)$, for sufficiently large $N$, a single iteration of the RPA decoder, which includes a run of an ML decoder for a first-order RM code and an aggregation step, produces the correct codeword as the decoded estimate.
For second-order RM codes, one can in addition obtain a bound on the error probability after two iterations of the RPA decoder, thereby yielding an improvement over the bound in Corollary IV.1. The following lemma makes this precise:
Lemma IV.5.
After two iterations of the RPA decoder for second-order RM codes, we have for all $\epsilon<\eta(p)$ that
$\Pr[\overline{\mathbf{Y}}=\mathbf{0}]\geq 1-256N^{2}\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
We thus incur a saving of a multiplicative factor that grows linearly in $N$, as compared to the bound in Corollary IV.1. We now prove Lemma IV.5.
Proof.
From Proposition IV.2, for any $\mathbf{x}\in\{0,1\}^{m}$, we have that $\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=Y_{\mathbf{x}}\ \bigg{|}\ \mathcal{G}\right]\geq 1-16N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ Via Markov’s inequality, this then allows us to obtain an upper bound on the probability that, conditioned on the event $\mathcal{G}$, the vector $\textsf{Flip}^{(N)}$ differs from $\mathbf{Y}$ in at least $2^{m-3}$ positions, where $22^{m-3}=N/8=d_{\text{min}}(\text{RM}(m,2))/$ 2. Indeed, we have that
$\Pr\left[w\left(\textsf{Flip}^{(N)}\oplus\mathbf{Y}\right)\geq N/8\mid\mathcal{G}\right]$ $\leq\frac{16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}}{N/8}$
$=128N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
By calculations similar to those in (8)–(9), we obtain $\Pr\left[w\left(\textsf{Flip}^{(N)}\oplus\mathbf{Y}\right)\geq N/8\right]\geq 1-256N^{2}e^{-\frac{N\epsilon^{2}}{8}}$. Now, note that if $\textsf{Flip}^{(N)}$ and $\mathbf{Y}$ differ in at most $N/8-1$ positions, we have that at the end of one RPA iteration, $w\left(\overline{\mathbf{Y}}\right)\leq d_{\text{min}}(\text{RM}(m,2))/2-1$. This hence implies, via arguments as in [25, Lemma 1] that the RPA decoder, in the next iteration, can correct all errors present in $\overline{\mathbf{Y}}$, resulting in convergence to the correct all-zeros codeword, at the end of two iterations. ∎
While Lemma IV.5 gives rise to a tighter estimate of the error probability of decoding second-order RM codes under RPA decoding, in this work, we persist with the calculations made via Theorem IV.2, since the above reduction in error probabilities leaves the exponent unchanged. Further, as we shall see in the next section, Theorem IV.2 can be used directly as an estimate of the error probabilities in the aggregation phase, for RM codes of order $r$ larger than $2$ as well.
V Extension to Higher-Order RM Codes
Via our analysis in the previous section for second-order RM codes, we obtained estimates (lower bounds) on the probability of correct decoding of the ML decoder for first-order RM codes, and the probability with which a single aggregation step produces the correct estimate $\mathbf{0}$ of a codeword, given that all of the projections of the received vector are decoded to $\mathbf{0}$. These results, as we argue in this section, allow us to obtain estimates of the probability of correct decoding under RPA decoding of general Reed-Muller codes RM$(m,r)$ too, thereby resulting in a proof of Theorem III.1.
First, we introduce some notation. Fix a Reed-Muller code RM$(m,r)$, for some $m\geq 1$ and $2<r\leq m$. Now, for any $t\in[r]$, let $\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(t)}$ be a sequence of one-dimensional subspaces, where $\mathbb{S}_{j}\subseteq\{0,1\}^{m-j+1}$, for each $j\in[t]$. Given the (random) received vector $\mathbf{Y}$ obtained by passing the input codeword $\mathbf{0}$ through the BSC$(p)$, we recall from (1) that
$\mathbf{Y}_{/\mathbb{S}^{(1)}}:=\left(\mathbf{Y}_{/\mathbb{S}^{(1)}}(T):\ T\in\{0,1\}^{m}/\mathbb{S}^{(1)}\right).$
The following quantities are then defined recursively: for every $11\leq j\leq t-$ 1, we index the coordinates of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}$ by vectors $\mathbf{z}\in\{0,1\}^{m-j}$ in lexicographic order, and set
$(\displaystyle\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j+1)}\right)}:=\bigg{(}$$\left(\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}\right)_{/\mathbb{S}^{(j+1)}}(T):$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ T\in\{0,1\}^{m-j}/\mathbb{S}^{(j+1)}\bigg{)}.$ (10)
For any (random) codeword $\mathbf{C}\in\text{RM}(m,r)$, one can define projections $\mathbf{C}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}$, for $1\leq j\leq m,$ analogously; via arguments made in [22, Lemma 1], it can be shown that $\mathbf{C}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}\in\text{RM}(m-j,r-j)$, if $\mathbf{C}\in\text{RM}(m,r)$.
The definitions above, along with the recursive structure of the RPA decoder (see Algorithm 1) motivate the definition of a “projection-aggregation tree"—a simple, visual device that helps us recursively compute the probabilities of error in decoding for higher-order RM codes. Formally, a projection-aggregation tree is a rooted tree, with the root (at level 00) being the code RM$(m,r)$. For every level $30\leq i\leq r-$ 3, each node at level $i$ has $\frac{N}{2^{i}}-1$ children at level $i+1$, each of which is an RM$(m-i-1,r-i-1)$ code; here, $N:=2^{m}$. Clearly, the leaves of the projection-aggregation tree are RM$(m-r+2,2)$ codes. Figure 1 shows a pictorial depiction of a projection-aggregation tree. Each path in the tree from the root to a leaf node corresponds to a fixed sequence $\mathbb{S}_{1},\ldots,\mathbb{S}_{r-2}$ of one-dimensional subspaces as above. “Moving down" from a parent node to a child node in this tree signifies a projection step and “moving up" from a child node to its parent signifies an aggregation step.
Now, for any $j\in[r-2]$, let $p^{(j)}:=\frac{1}{2}\cdot(1-(1-2p)^{2^{j}})$ and $N^{(j)}:=N/2^{j}$. Since $\mathbf{Y}\sim\text{Ber}^{\otimes N}(p)$, the following simple lemma holds:
Lemma V.1.
For any choice of one-dimensional subspaces $\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}$, for $j\in[r-2]$, where $\mathbb{S}_{k}\subseteq\{0,1\}^{m-k+1}$, for $k\in[j]$, we have
$\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}\sim\text{Ber}^{\otimes(N^{(j)})}(p^{(j)}).$
Proof.
First, observe that the vector $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ is of length $N/2^{j}=N^{(j)}$, by construction. Further, by the disjointness of distinct cosets, we obtain from (V) that each coordinate of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ is the sum, modulo $2$, of $2^{j}$ distinct coordinates of $\mathbf{Y}$. This implies that any coordinate of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ takes the value $1$ with probability equal to the probability that the parity, or the sum modulo $2$ of $2^{j}$ Ber$(p)$ random variables, equals $1$, which in turn equals $p^{(j)}$. Moreover, distinct coordinates of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ correspond to modulo $2$ sums of disjoint collections of coordinates in $\mathbf{Y}$, thereby yielding the statement of the lemma. ∎
Now, consider any leaf node in the projection-aggregation tree, corresponding to a second-order RM code RM$(m-r+2,2)$, obtained via some fixed path of subspaces $\overline{\mathbb{S}}_{1},\ldots,\overline{\mathbb{S}}_{r-2}$. Let $\textsc{Agg}^{(r-2)}\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)}\right)$ denote the vector $\overline{\mathbf{Y}}$ obtained at the node of interest, post aggregation using the decoded estimates of the projections
$\left(\mathbf{Y}_{/\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)},\mathbb{S}^{(r-1)}\right)}:\ \mathbb{S}^{(r-1)}\in\{0,1\}^{m-r+2}\right)$
Recall that the projections above were decoded using the ML decoder for first-order RM codes. Using the fact (from Lemma V.1) that $\mathbf{Y}_{/\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)}\right)}\sim\text{Ber}^{\otimes(N^{(r-2)})}(p^{(r-2)})$, we obtain via our analysis of the error probabilities for second-order RM codes (in particular, Corollary IV.1) that for any $\epsilon<\eta(p^{(r-2)})$,
$\Pr$ $\left[\textsc{Agg}^{(r-2)}\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)}\right)=\mathbf{0}\right]\geq$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 1-32\left(N^{(r-2)}\right)^{3}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}.$
Now, consider the (unique) parent of the child node of interest, which corresponds to the path $\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)}$. Observe that the vectors $\textsc{Agg}^{(r-2)}\left(\overline{\mathbb{S}}^{(1)},\ldots,{\mathbb{S}}^{(r-2)}\right)$ are precisely the $\widehat{\mathbf{Y}}_{/{\mathbb{S}}^{(r-2)}}$ vectors that are passed “upwards" in the projection-aggregation tree to the parent node (see Step 6 in Algorithm 1). For notational consistency, we denote each of these “aggregated" vectors as $\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}$. Via a simple union bound, we see that the probability that each of $\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}$, corresponding to each of the children of the parent node under consideration, equals $\mathbf{0}$, obeys
$\Pr$ $\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}=\mathbf{0},\ \text{for all ${\mathbb{S}}^{(r-2)}$}\right]\geq$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 1-32\left(N^{(r-2)}\right)^{4}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}},$ (11)
for any $\epsilon<\eta(p^{(r-2)})$. Thus, defining the event
$\mathcal{G}^{(r-2)}:=\left\{\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}=\mathbf{0},\ \text{for all ${\mathbb{S}}^{(r-2)}$}\right\},$
we observe that the result in (11) above is entirely analogous to Proposition IV.1.
Similar to the definitions of the “aggregated" vector $\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}$, one can define, for every $j\in[r-3]$, the vector $\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(j-1)},\mathbb{S}^{(j)}}\right)}$ corresponding to the children of the node indexed by the path $\overline{\mathbb{S}}^{(1)},\ldots,\mathbb{S}^{(j-1)}$. Likewise, for any $j\in[r-3]$, we define the event
$\mathcal{G}^{(j)}:=\left\{\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(j-1)},\mathbb{S}^{(j)}}\right)}=\mathbf{0},\ \text{for all ${\mathbb{S}}^{(j)}$}\right\}.$
The following theorem then holds:
Theorem V.1.
For any $j\in[r-2]$, we have that for any $\epsilon<\eta(p^{(r-2)})$,
$\Pr\left[\mathcal{G}^{(j)}\right]\geq 1-2^{r+3-j}\cdot\left(N^{(j)}\right)^{r+2-j}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}.$
Proof.
We proceed by induction on $j$. From (11), we see that the statement of the theorem holds for $j=r-2$. Now, assume that the statement holds for some $i<r-2$. Then, arguing similar to the sequence of inequalities (8)–(9) in the proof of Theorem IV.2 (and implicitly using Proposition IV.2), we see that for any subspace $\mathbb{S}^{(i-1)}\subseteq\{0,1\}^{i-1}$,
$\Pr\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(i-2)},\mathbb{S}^{(i-1)}}\right)}=\mathbf{0}\right]$
$\geq 1-16\left(N^{(i-1)}\right)^{2}\left(N^{(i-1)}-1\right)\cdot e^{-\frac{N^{(i-1)}\epsilon^{2}}{8}}-$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 2^{r+3-i}\cdot\left(N^{(i)}\right)^{r+2-i}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$
$\geq 1-16\left(N^{(i-1)}\right)^{3}\cdot e^{-\frac{N^{(i-1)}\epsilon^{2}}{8}}-$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 2\cdot\left(N^{(i-1)}\right)^{r+2-i}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$
$\geq 1-32\cdot\left(N^{(i-1)}\right)^{r+2-i}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$
$\geq 1-2^{r+4-i}\cdot\left(N^{(i-1)}\right)^{r+2-i}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$
where the third inequality makes use of the fact that $i<r-2\leq m-2$, implying that $N^{(i)}>4$. Therefore, by a union bound,
$\Pr\left[\mathcal{G}^{(i-1)}\right]$ $=\Pr\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(i-2)},\mathbb{S}^{(i-1)}}\right)}=\mathbf{0},\ \text{for all $\mathbb{S}^{(i-1)}$}\right]$
$\geq 1-2^{r+4-i}\cdot\left(N^{(i-1)}\right)^{r+3-i}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}},$
thereby proving the inductive step. ∎
The theorem above immediately results in a proof of Theorem III.1:
Proof of Theorem III.1.
From Theorem V.1, we see that for any $\epsilon<\eta(p^{(r-2)})$,
$\Pr\left[\mathcal{G}^{(1)}\right]$ $=\Pr\left[\widehat{\mathbf{Y}}_{/{{\mathbb{S}}^{(1)}}}=\mathbf{0},\ \text{for all $\mathbb{S}^{(1)}$}\right]$
$\geq 1-2^{r+2}\cdot(N/2)^{r+1}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$
$=1-2N^{r+1}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}.$
Thus, we obtain, via analysis similar to that in the set of inequalities (8)–(9) in the proof of Theorem IV.2, that
$P_{\text{err}}(\text{RM}(m,r))$
$\leq 16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}+2N^{r+1}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$
$\leq 32N^{r+1}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}=32N^{r+1}\cdot e^{-2^{-(r+1)}\cdot{N\epsilon^{2}}},$
thereby proving our main result. ∎
We end this section with a remark. While our analysis of the performance of RPA decoding has restricted attention to projections via one-dimensional subspaces, it will be interesting to study precise estimates of the probabilities of error under an RPA decoding procedure that makes use of higher dimensional subspaces. Based on analysis done in this paper, we believe that the use of higher dimensional subspaces for projection will indeed lead to tighter bounds on $P_{\text{err}}(\text{RM}(m,r))$, as against that in Theorem III.1, even when we restrict attention to the event that the RPA decoder converges after a single iteration. Our intuition for why this may be true is that the union bound arguments made in our analysis appear to scale the (exponentially small) probability of error of decoding a first-order RM code by a factor that roughly equals the total number of edges in the projection-aggregation tree. If one were to use higher dimensional subspaces for projection, the total number of edges in the corresponding projection-aggregation tree can be made significantly smaller. We leave a precise analysis of the probabilities of error in such cases for future work.
VI Conclusion and Future Work
In this paper, we provided a formal, theoretical analysis of the performance of the Recursive Projection-Aggregation (RPA) decoder for Reed-Muller (RM) codes over a binary symmetric channel (BSC). Via our analysis, we obtained that the RPA decoder recovers the correct input codeword with overwhelming probability, for RM codes of order logarithmic in $m$, for large enough blocklengths $N=2^{m}$. Key components of our analysis were our estimates of the error probabilities of the Fast Hadamard Transform-based decoder and the aggregation step in the RPA decoder.
An interesting line of future work would be to extend the analysis in this paper to situations where more than one iteration of the RPA decoder is run, before convergence. We believe that such an attempt will allow one to obtain vanishing error probabilities for higher RM orders than guaranteed in this paper. Another important extension of our analysis will be to the setting of the transmission of RM codewords over general binary-input, memoryless, symmetric (BIMS) channels (or even over general discrete memoryless channels or DMCs), of which the BSC is a special case.
Acknowledgement
This work was carried out when the second author was visiting the Indian Institute of Science, Bengaluru. The authors thank Prof. Navin Kashyap and Prof. Henry Pfister for helpful comments and suggestions.